OpenStudy (anonymous):

m=2 intersects 2x-3y=6 at x=3 what is the slope intercept form?

OpenStudy (anonymous):

we know if x = 3 then \[6-3y=6\] \[-3y=0\] \[y=0\] so we have the point (3,0) and the slope 2 so \[y-0=2(x-3)\] \[y=2x-6\]